OpenStudy (anonymous):

Bob’s Boat Rentals charges a flat fee of $80 plus $5/h to rent a boat. Larry’s Boat Rentals charges $15/h with no flat fee. The system that models this situation is given, where c is the cost of renting a boat and h is the number of hours a boat is rented. The solution to the system is (8, 120). What is the correct interpretation of the solution of this system? c=80+5h c=15h

OpenStudy (anonymous):

A. The cost to rent a boat is the same with both boat rental companies if you rent a boat for 8 h. Both companies will charge $120. B. The cost to rent a boat is the same with both boat rental companies if you rent a boat for 120 h. Both companies will charge $15. C. Bob’s Boat Rental will charge $120 for a rental of 8 hours. D. Bob’s Boat Rental will charge less money on the fifth day, charging $140.

ganeshie8 (ganeshie8):

what do you think ?

OpenStudy (anonymous):

my thought would be A

OpenStudy (texaschic101):

lets check.. Bob's boat for 8 hrs.. c = 80 + 5(8) = 80 + 40 = 120 Larry's boat for 8 hrs.. c = 15(8) = 120 yep....I agree with you. I believe it is A.

OpenStudy (anonymous):

thanks

OpenStudy (texaschic101):

sure thing :)